OpenStudy (anonymous):

Topic: Sinking Fund Factor [Just need an explanation] Jim Hayes wants to buy some electronic equipment for $1000. He decided to save a uniform amount at the end of each month so that he would have the required $1000 at the end of one year. The local credit union pays 6% interest, compounded monthly. How much would Jim have to deposit each month?

OpenStudy (anonymous):

Solution: F= $ 1, 000 n= 12 i= 0.5% A= unknown A= F ( A/F, i%, n ) A= 1000(0.0811) = $81.10 [Note: ( A/F, 0.5%, 12 ) value came from a compound interest table] \(\sf My\ Question:\) Why is i=0.5% not 6%?

Miracrown (miracrown):

That's because the questions tells us the interest rate is 6%, which means it's 6% for the whole year, but they say compounded monthly so that means 6%/12 which is 0.5%. They told us compounded monthly, so it's not 6% per month, it's 6%/12 because there are 12 months in a year.

OpenStudy (anonymous):

oh that's it /.- I must be sleepy now lol Thank you so much! :D

Miracrown (miracrown):

Hehe. Get some rest! And no worries. :^)

OpenStudy (anonymous):

will do, goodnight :)

Miracrown (miracrown):

night, sleep tight